Combine like terms.
x + 6y + 15 - 7x - 3y - 13

Answer by fractalier(2101)

From x + 6y + 15 - 7x - 3y - 13, we add all of the x's together, all of the y's together, and all of the number (or constant) terms together. Thus,
x + 6y + 15 - 7x - 3y - 13 =
(x - 7x) + (6y - 3y) + (15 - 13) =
-6x + 3y + 2